OpenStudy (anonymous):

Solve 2x^2+20x+3 =0 by completing the square

sam (.sam.):

\[2x^2+20x+3 =0\] \[2(x^2+10x)+3=0\] \[2(x+5)^2-25(2)+3=0\] \[2(x+5)^2-47=0\] Solve for x

OpenStudy (anonymous):

thanks, but why isn't 3 divided by two when you factorise?

sam (.sam.):

There's a few techniques to solve this

sam (.sam.):

which grade you in?

OpenStudy (anonymous):

15 years old, not sure educational system in US but i think i'm in grade 10

sam (.sam.):

My technique is, \[ax^2+bx+c=0\] Always make sure that "a" is 1 Factor out the coefficient of x^2 \[a(x^2+\frac{b}{a}x)+c=0\] Then, divide 2 from b/a , then bring out b/2a (always minus sign) and square b/2a , then stick back "a" to b/2a \[a(x+\frac{b}{2a})^2-(\frac{b}{2a})^2(a)+c=0\]

OpenStudy (anonymous):

thank you, helped a lot with my revision :)

sam (.sam.):

its quite confusing at first but if you get the hang of it, its the fastest method to solve a completing the quare